40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 16975     Entry time: Wed Jul 6 19:58:16 2022
Author: Paco 
Type: Summary 
Category: NoiseBudget 
Subject: XARM noise budget 

[Anchal, Paco, Rana]

We locked the XARM using POX11 and made a noise budget for the single arm displacement; see Attachment #1. The noise budget is rough in that we use simple calibrations to get it going; for example we calibrate the measured error point C1:LSC-XARM_IN1_DQ using the single cavity pole and some dc gain to match the UGF point. The control point C1:LSC-XARM_OUT_DQ is calibrated using the actuator gain measured recently by Yuta. We also overlay an estimate of the seismic motion using C1:PEM-SEIS_BS_X_OUT_DQ (calibrated using a few poles to account for stack and pendulum), and finally the laser frequency noise as proxied by the mode cleaner C1:IOO-MC_F_DQ.

A couple of points are taken with this noise budget, apart from it needing a better calibration;

  1. Overall the inferred residual displacement noise is high, even for our single arm cavity.
    1. By looking at the sim OLTF in foton, it seemed that the single arm cavity loop TF could easily become unstable due to some near-UGF-funkiness likely from FM3 (higher freq boost), so we disabled the automatic triggering on it; the arm stayed locked and we changed the error signal (light blue vs gold (REF1) trace)
  2. The arm cavity is potentially seeing too much noise from the IMC in the 1 to 30 Hz band in the form of laser frequency noise.
    1. Need IMC noise budget to properly debug.
  3. At high frequency (>UGF), there seem to be a bunch of "wiggles" which remain unidentified.
    1. We actually tried to investigate a bit into these features, thinking they might have something to do with misalignment, but we couldn't really find significant correlation.

RXA edit:

  1. we also noticed some weirdness in the calibration of MC_F v. Arm. We think MC_F should be in units of Hz, and Paco calculated the resulting motion as seen by the arm, but there was a factor of several between these two. Need to calibrate MC_F and check. In principle, MC_F will show up directly in ALS_BEATX (with the green PDH lock off), and I assume that one is accurately calibrated. Somehow we should get MC_F, XARM, and ALS_BEAT to all agree. JC is working on calibrating the Mini-Circuits frequency counter, so once that is done we will be in good shape.
  2. we may need to turn on some MC_L feedback for the IMC, so that the MC length follows the NPRO frequency below ~20 Hz.
  3. Need to estimate where the IMC WFS noise is in all of this. Does it limit the MC length stability in any frequency band? How do we determine this?
  4. Also, we want to redo this noise budget today, whilst using AS55 instead of POX. Please measure the Schnupp asymmetry by checking the optimum demod phase in AS55 for locking Xarm v Yarm.
Attachment 1: xarm_nb_2022_07.pdf  62 kB  Uploaded Wed Jul 6 21:47:56 2022  | Hide | Hide all
xarm_nb_2022_07.pdf
ELOG V3.1.3-